# Adapting Vision Foundation Models for Robust Cloud Segmentation in Remote Sensing Images

### v1

@@ Abstract
Cloud segmentation is a critical challenge in remote sensing image interpretation, as its accuracy directly impacts the effectiveness of subsequent data processing and analysis. Recently, vision foundation models (VFM) have demonstrated powerful generalization capabilities across various visual tasks. In this paper, we present a parameter-efficient adaptive approach, termed Cloud-Adapter, designed to enhance the accuracy and robustness of cloud segmentation. Our method leverages a VFM pretrained on general domain data, which remains frozen, eliminating the need for additional training. Cloud-Adapter incorporates a lightweight spatial perception module that initially utilizes a convolutional neural network (ConvNet) to extract dense spatial representations. These multi-scale features are then aggregated and serve as contextual inputs to an adapting module, which modulates the frozen transformer layers within the VFM. Experimental results demonstrate that the Cloud-Adapter approach, utilizing only 0.6% of the trainable parameters of the frozen backbone, achieves substantial performance gains. Cloud-Adapter consistently attains state-of-the-art (SOTA) performance across a wide variety of cloud segmentation datasets from multiple satellite sources, sensor series, data processing levels, land cover scenarios, and annotation granularities. We have released the source code and pretrained models at https://github.com/XavierJiezou/Cloud-Adapter to support further research.
@@ I Introduction
Figure 1: Comparison between (a) the Previous Method and (b) our Cloud-Adapter approach. Unlike (a), where the entire network (e.g., a typical encoder-decoder architecture) is fully trainable, resulting in a large number of parameters and an increased risk of overfitting, (b) our Cloud-Adapter approach leverages a frozen vision foundation model (VFM) combined with a lightweight, trainable adapter. This design preserves generalization and adaptability, enabling efficient learning for cloud segmentation tasks. The frozen VFM extracts generic visual features, while the parameter-efficient adapter facilitates effective transfer learning across remote sensing scenes.
Cloud segmentation, also known as cloud detection, plays a critical role in the interpretation of remote sensing images, with wide applications in climate monitoring [1, 2], environmental protection, disaster warning, and land use [3]. The variations and characteristics of clouds significantly influence the quality of remote sensing images and subsequent analysis results. Therefore, accurate cloud segmentation is essential for extracting meaningful geospatial information and ensuring reliable data processing. However, due to the diversity of cloud shapes, textures, and sizes, coupled with the complexity of imaging conditions, traditional cloud segmentation methods [4, 5, 6, 7] often fail to achieve stable high-accuracy results in dynamic remote sensing images. Consequently, developing a more robust and adaptable cloud segmentation method has become a key research focus in the field of remote sensing.
With the rapid development of deep learning [8], convolutional neural network (ConvNet [9, 10, 11])-based cloud segmentation methods [12, 13, 14, 15, 16, 17, 18] have made remarkable progress, particularly due to their ability to automatically extract high-level features that enhance the accuracy of cloud shape, edge detection, and texture analysis. For instance, MFCNN [14] combines multi-scale convolutional features to enable fine-grained cloud detection, while DABNet [18] introduces the deformable context feature pyramid module to improve the adaptive modeling capability of multi-scale features for high-accuracy cloud detection. Despite these advancements, many methods still face challenges, including strong dependence on training data, high model complexity, and limited adaptability to diverse remote sensing scenes.
@@
To address these challenges, we propose a novel cloud segmentation method based on vision foundation models, named Cloud-Adapter (see Fig. 1). This approach freezes the pretrained VFM backbone and incorporates a lightweight adaptive module, requiring minimal additional parameters to enhance cloud segmentation accuracy and robustness. Specifically, we design a spatial perception module that employs convolutional neural networks (ConvNets) to extract dense spatial features and aggregate them into multi-scale contextual information. This information is then passed into the adaptive module, which modulates the frozen transformer layers of the VFM to efficiently perform cloud segmentation. Experimental results demonstrate that Cloud-Adapter achieves significant performance improvements with only 0.6% of the trainable parameters of the backbone across various popular datasets.
@@
    We demonstrate that our method consistently achieves state-of-the-art performance on four datasets (six subsets) including multiple satellite sources, sensor types, and land cover scenarios, highlighting its robustness towards remote sensing cloud segmentation tasks.
@@ III Method
In this section, we introduce a simple but powerful fine-tuning method called Cloud-Adapter (as shown in Fig. 2) for cloud segmentation by freezing the vision foundation model (VFM) and incorporating a lightweight adapting module. The method consists of two key components: spatial perception and adapting modules. The spatial perception module extracts multi-scale spatial features using multiple ConvNet blocks and then aggregates them into a dense context. The adapting modules modulate the frozen VFM layers with the context, enabling effective transfer learning for cloud segmentation.
@@ III-A Spatial Perception Module
After the stem, the input features are processed by multiple ConvNet blocks, each of which extracts spatial features at different scales, resulting in a series of feature maps $\mathbf{F}_{i}^{c}\in\mathbb{R}^{C_{i}\times H\times W}$, where $i$ denotes different scales. To aggregate these multi-scale feature maps into a unified scale, we use an aggregator, which applies adaptive average pooling to normalize the feature maps to the same size. The aggregator is a parameter-free operation, meaning it does not contain any trainable parameters, making it a zero-parameter operation.
@@
$\mathbf{F}_{\text{agg}}=\sum_{i=1}^{k}\mathcal{P}_{avg}(\mathbf{F}_{i}^{c},\text{output\_size}=\mathbf{F}_{k}^{c}.\text{shape}[-2:]),$ (1)
where $\mathbf{F}_{i}^{c}$ are the feature maps at different scales, and $\mathcal{P}_{avg}$ is the adaptive average pooling operation that pools the feature maps to the smallest size of the final feature map $\mathbf{F}_{k}^{c}$.
@@ III-B Adapting Module
The adapting module, as referred to in Fig. 3, utilizes the cross-attention mechanisms [52, 53, 54, 55] to adjust the frozen VFM’s Transformer layers. Specifically, it takes the context $\mathbf{F}_{\text{agg}}$ from the spatial perception module and interacts with the frozen VFM features, thereby modulating the output of the VFM. The cross-attention mechanism calculates the relationships between aggregated dense features and the features extracted by the VFM, producing the adapted features.
The cross-attention computation is expressed by:
$S=\text{Softmax}\left(\frac{\mathbf{QK}^{T}}{\sqrt{d}}\right),$ (2)
@@
where $\mathbf{Q}\in\mathbb{R}^{d}$, $\mathbf{K}\in\mathbb{R}^{d}$, and $\mathbf{V}\in\mathbb{R}^{d}$ are the query, key, and value matrices, respectively. The Softmax operation computes the attention weights $\mathbf{S}$, which are then multiplied with the value matrix $\mathbf{V}$ to obtain the output $\mathbf{O}$. Finally, after the mapping operation (denoted as $\mathcal{M}$), the adapted features $\mathbf{F}_{j}^{att}$ are obtained with a residual connection, where $j$ denotes the $j$-th feature from the frozen VFM transformer layers.
@@ IV-A4 HRC_WHU
[12] HRC_WHU is a high-resolution cloud cover validation dataset created by the SENDIMAGE Lab at Wuhan University. It comprises 150 high-resolution images acquired from Google Earth with RGB channels and resolutions varying from 0.5 to 15 meters. These images have been expertly digitized with two mask annotations: clear sky and clouds. Following the original paper, We uniform cut to 256 $\times$ 256 patch size, 120 for training, and 30 for testing.
@@ IV-B Implementation Details
To ensure experimental reproducibility, this section provides a comprehensive overview of the hyperparameter configurations used during training and the evaluation metrics applied to assess the experimental results. All experiments were conducted on a workstation equipped with NVIDIA GeForce RTX 3090 GPUs with 24 GB of memory. The operating system was Ubuntu 22.04, and the neural networks were implemented in PyTorch 2.4.1 with CUDA 11.8. The experiments were performed within the MMSegmentation [58] framework (version 1.2.2), a sound semantic segmentation toolbox based on PyTorch. This framework offers an intuitive interface and a variety of pre-built models, facilitating streamlined experimentation. All segmentation heads across the models used in this study were based on Mask2Former [59], a flexible and efficient architecture tailored for image segmentation tasks. Mask2Former’s unified approach to segmentation and detection tasks contributed to consistent and reliable performance across the different models evaluated in these experiments.
@@ IV-B1 Training Settings
We employed a custom learning rate schedule to optimize model performance. The schedule began with a warm-up phase, utilizing a linear scheduler that increased the learning rate from $1\times 10^{-6}$ over the first 1,000 iterations. This was followed by a PolyLR scheduler with an eta minimum of 0.0 and a power of 0.9, applied from iteration 1,000 to 40,000. The training loop was configured to run for a maximum of 40,000 iterations, with validation conducted at intervals of every 4,000 steps. To ensure reproducibility, we set the random seed to 42 and used the AdamW optimizer [60] with an initial learning rate of $1\times 10^{-4}$ and a weight decay of 0.05. The batch size during training was set to 4, and parameters were initialized using the kaiming normal [61].
@@ IV-B2 Evaluation Metrics
The metrics mAcc, mIoU, and mDice represent the mean values computed for each class, offering insight into the model’s performance at the class level. In contrast, the overall accuracy (aAcc) evaluates the model’s classification performance across the entire dataset. However, aAcc often overlooks minority classes, such as cloud shadow and thin cloud, potentially overestimating the model’s effectiveness. A significant discrepancy between aAcc and mAcc may highlight class imbalance, where dominant classes skew results. To address this, we calculate the average of these metrics across all courses for cloud segmentation, ensuring a balanced evaluation by considering both per-class accuracy and overall pixel-wise accuracy. These metrics are all positive indicators, with higher values signifying better performance, simplifying interpretation, and facilitating model comparison.
@@ IV-C Ablation Studies and Analysis
This section presents a series of comprehensive ablation studies to evaluate the effectiveness of key components in our proposed model using the CloudSEN12_High_L1C dataset. Specifically, we investigate the impact of (1) different variants of the VFM backbone, (2) the ConvNet architectures employed in the SPM, (3) the dimensions of interactive context, (4) the low-rank MLP in the adapting module, (5) the importance of the aggregator, (6) the interaction frequency within the adapting module, and (7) the presence or absence of each component in the Cloud-Adapter. These studies provide a detailed analysis of how each component contributes to cloud segmentation performance, ultimately highlighting the strengths and advantages of our proposed Cloud-Adapter.
@@ IV-C1 Variants of VFM’s Backbone
To evaluate the impact of different vision foundation models and their variants on our method’s performance, we conducted comprehensive experiments using SAM [22] and DINOv2 [23] as backbones. Table II compares three available model variants of them.
We observed that as the model scale increased, both SAM and DINOv2 exhibited consistent performance improvements. Among the SAM variants, the Huge version performed the best, achieving 73.70% mIoU, 84.37% mAcc, 89.95% aAcc, and 84.07% mDice. In contrast, DINOv2 (Large) demonstrated superior performance across all metrics, reaching 74.18% mIoU, 84.79% mAcc, 90.19% aAcc, and 84.46% mDice.
@@
These results indicate that DINOv2 (Large) outperforms SAM (Huge) across all metrics while being significantly more efficient in terms of training time and memory usage. Therefore, DINOv2 (Large) emerges as the optimal backbone for our method, striking an excellent balance between performance and computational efficiency for cloud segmentation.
@@ IV-C3 Dimension of Interactive Context
The results show that the model’s performance does not necessarily improve with larger dimensions. Starting from dimension 16 with a mIoU of 73.72%, we observe a slight improvement to 73.77% when increasing to dimension 32. The peak performance is achieved at dimension 64 with a mIoU of 74.18%. However, further increasing the dimension to 128 leads to a slight performance degradation (0.44% dropping).
@@ IV-C4 Low-Rank MLP of Adapting Module
The experimental results reveal an interesting relationship between rank size and model performance. With a rank of 4, the model achieves a mIoU of 73.84% while requiring only 0.61M trainable parameters. Increasing the rank to 8 improves performance to 74.03% with 1.02M parameters. The peak performance is reached at rank 16 with a mIoU of 74.18%, which requires 1.82M parameters. Further increasing the rank to 32 leads to a slight performance degradation. Based on these observations, we choose rank 16 as our optimal setting. This finding aligns with LoRA’s design principle [46] of achieving efficient adaptation through low-rank approximation.
@@ IV-C5 Importance of the Aggregator
The “Aggregated” strategy, which fuses multiple scale features into a single output with parameter-free operation, demonstrates the best overall performance. It achieves 74.18% mIoU, 84.79% mAcc, and 84.46% mDice, while maintaining the same aAcc (90.19%) as the “Single-Scale” approach. This superior performance can be attributed to the strategy’s ability to effectively combine and utilize information from different feature scales. Based on these results, we adopt the “Aggregated” feature output strategy in our final model design, as it provides the most effective use of dense multi-scale feature information without introducing additional parameters.
@@ IV-C6 Interaction Frequency of Adapting Module
The results indicate that the optimal performance is achieved when $N$ is set to 24, where the model attains 74.18% mIoU, 84.79% mAcc, 90.19% aAcc, and 84.46% mDice. This configuration outperforms all others, particularly in terms of mIoU and mAcc. Notably, the increase in interaction frequency from 4 to 24 leads to improvements in the model’s ability to capture complex spatial dependencies across layers, significantly boosting the performance metrics.

### v2

@@ Abstract
Cloud segmentation is a critical challenge in remote sensing image interpretation, as its accuracy directly impacts the effectiveness of subsequent data processing and analysis. Recently, vision foundation models (VFM) have demonstrated powerful generalization capabilities across various visual tasks. In this paper, we present a parameter-efficient adaptive approach, termed Cloud-Adapter, designed to enhance the accuracy and robustness of cloud segmentation. Our method leverages a VFM pretrained on general domain data, which remains frozen, eliminating the need for additional training. Cloud-Adapter incorporates a lightweight spatial perception module that initially utilizes a convolutional neural network (ConvNet) to extract dense spatial representations. These multi-scale features are then aggregated and serve as contextual inputs to an adapting module, which modulates the frozen transformer layers within the VFM. Experimental results demonstrate that the Cloud-Adapter approach, utilizing only 0.6% of the trainable parameters of the frozen backbone, achieves substantial performance gains. Cloud-Adapter consistently achieves state-of-the-art performance across various cloud segmentation datasets from multiple satellite sources, sensor series, data processing levels, land cover scenarios, and annotation granularities. We have released the code and model checkpoints at https://xavierjiezou.github.io/Cloud-Adapter/ to support further research.
@@ I Introduction
Figure 1: Comparison between (a) the Previous Method and (b) our Cloud-Adapter approach. Unlike (a), where the entire network (e.g., a typical encoder-decoder architecture) is fully trainable, resulting in a large number of parameters and an increased risk of overfitting, (b) our Cloud-Adapter approach leverages a frozen vision foundation model (VFM) combined with a lightweight, trainable adapter. This design preserves generalization and adaptability, enabling efficient learning for cloud segmentation tasks. The frozen VFM extracts generic visual features, while the parameter-efficient adapter facilitates effective transfer learning on diverse remote sensing scenes.
Cloud segmentation, also known as cloud detection, plays a critical role in the interpretation of remote sensing images, with wide applications in climate monitoring [1, 2], environmental protection, disaster warning, and land use [3]. The variations and characteristics of clouds significantly influence the quality of remote sensing images and subsequent analysis results. Therefore, accurate cloud segmentation is essential for extracting meaningful geospatial information and ensuring reliable data processing. However, due to the diversity of cloud shapes, textures, and sizes, coupled with the complexity of imaging conditions, traditional cloud segmentation methods [4, 5, 6, 7] often fail to achieve stable high-accuracy results in dynamic remote sensing images. Consequently, developing a more robust and adaptable cloud segmentation method has become a key research focus in remote sensing.
With the rapid development of deep learning [8, 9, 10, 11], convolutional neural network (ConvNet)-based cloud segmentation methods [12, 13, 14, 15, 16, 17, 18] have made remarkable progress, particularly due to their ability to automatically extract high-level features that enhance the accuracy of cloud shape, edge detection, and texture analysis. For instance, MFCNN [14] combines multi-scale convolutional features to enable fine-grained cloud detection. At the same time, DABNet [18] introduces the deformable context feature pyramid module to improve the adaptive modeling capability of multi-scale features for high-accuracy cloud detection. Despite these advancements, many methods still face challenges, including strong dependence on training data, low model generalization, and limited adaptability to diverse remote sensing scenes.
@@
To address these challenges, we propose a novel cloud segmentation method based on vision foundation models named Cloud-Adapter (see Fig. 1). This approach freezes the pretrained VFM backbone and incorporates a lightweight adaptive module, requiring minimal additional parameters to enhance cloud segmentation accuracy and robustness. Specifically, we design a spatial perception module that employs convolutional neural networks (ConvNets) to extract dense spatial features and aggregate them into multi-scale contextual information. This information is then passed into the adaptive module, which modulates the frozen transformer layers of the VFM to perform cloud segmentation efficiently. Experimental results demonstrate that Cloud-Adapter achieves significant performance improvements with only 0.6% of the trainable parameters of the backbone across various popular datasets.
@@
    We demonstrate that our method consistently achieves SOTA performance on four datasets (six subsets), including multiple satellite sources, sensor types, and land cover scenarios, highlighting its robustness towards remote sensing cloud segmentation tasks.
@@ III Method
In this section, we introduce a simple but powerful adaptive method called Cloud-Adapter (as shown in Fig. 2) to achieve fine-tuning of cloud segmentation by freezing the weights of the visual foundation model (VFM). The method consists of two key components: the spatial perception module and the adapting modules. The spatial perception module extracts multi-scale spatial features using multiple ConvNet blocks and then aggregates them into a dense context. The adapting modules modulate the frozen VFM layers with the context, enabling effective transfer learning for cloud segmentation.
@@ III-A Spatial Perception Module
After the stem, the input features are processed by multiple ConvNet blocks, each of which extracts spatial features at different scales, resulting in a series of feature maps $\mathbf{F}_{i}^{c}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$, where $i$ denotes different scales. To aggregate these multi-scale feature maps into a unified scale, we use an aggregator, which applies adaptive average pooling to normalize the feature maps to the same size. The aggregator is a parameter-free operation, meaning it does not contain any trainable parameters, making it a zero-parameter operation.
@@
$\mathbf{F}_{\text{agg}}=\sum_{i=1}^{k}\mathcal{P}_{avg}(\mathbf{F}_{i}^{c}),$ (1)
where $\mathbf{F}_{i}^{c}$ are the feature maps at different scales, and $\mathcal{P}_{avg}$ is the adaptive average pooling operation that pools the feature maps to the smallest size of the final feature map $\mathbf{F}_{k}^{c}$. Additionally, $k\in\{1,2,\dots,\lfloor\log_{2}(\min(H_{0},W_{0}))\rfloor\}$ refers to the number of ConvNet blocks, where each block has a convolution with the stride of 2, achieving a downsampling effect by squeezing the spatial dimensions at each step. Using the aggregator, we integrate multi-scale hierarchical representations into a single dense prior $\mathbf{F}_{\text{agg}}\in\mathbb{R}^{C_{k}\times H_{k}\times W_{k}}$.
@@ III-B Adapting Module
The adapting module, as referred to in Fig. 3, utilizes the cross-attention mechanisms [52, 53, 54, 55] to adjust the frozen VFM’s Transformer layers. Specifically, it takes the context $\mathbf{F}_{\text{agg}}$ from the spatial perception module and interacts with the frozen VFM features, thereby modulating the output of the VFM. The cross-attention mechanism calculates the relationships between aggregated dense features and the features extracted by the VFM, producing the adapted features. The queries, keys, and values are computed as follows:
$\mathbf{Q}=\mathcal{M}_{q}(\mathbf{F}_{j}^{t}),\quad\mathbf{K}=\mathcal{M}_{k}(\mathbf{F}_{\text{agg}}),\quad\mathbf{V}=\mathcal{M}_{v}(\mathbf{F}_{\text{agg}}),$ (2)
where $\mathbf{Q}\in\mathbb{R}^{d}$, $\mathbf{K}\in\mathbb{R}^{d}$, and $\mathbf{V}\in\mathbb{R}^{d}$ are the query, key, and value matrices, respectively. $\mathcal{M}_{q}$, $\mathcal{M}_{k}$, and $\mathcal{M}_{v}$ denote learnable linear projection functions for generating the respective matrices. $\mathbf{F}_{j}^{t}$ represents the robust feature extracted from the frozen VFM transformer layers, where $j$ indicates the index. Then cross-attention computation is expressed by:
$\mathbf{S}=\text{Softmax}\left(\frac{\mathbf{QK}^{T}}{\sqrt{d}}\right),$ (3)
@@
where the Softmax operation computes the attention weights $\mathbf{S}$, which are then multiplied with the value matrix $\mathbf{V}$ to obtain the output $\mathbf{O}$. Finally, after the mapping operation (denoted as $\mathcal{M}$), the adapted features $\mathbf{F}_{j}^{att}$ are obtained with a residual.
@@ IV-A4 HRC_WHU
[12] HRC_WHU is a high-resolution cloud cover validation dataset. It comprises 150 high-resolution images acquired from Google Earth with RGB channels and resolutions varying from 0.5 to 15 meters. These images have been expertly digitized with two mask annotations: clear sky and clouds. Here, We uniform cut to 256 $\times$ 256 patch size, 120 for training, and 30 for testing.
@@ IV-B1 Training Settings
All experiments were conducted on a workstation with NVIDIA GeForce RTX 3090 GPUs. The number of ConvNet blocks ($k$) was set to 4. We employed a custom learning rate schedule to optimize model performance. The schedule began with a warm-up phase, utilizing a linear scheduler that increased the learning rate from $1\times 10^{-6}$ over the first 1,000 iterations. This was followed by a PolyLR scheduler with an eta minimum of 0.0 and a power of 0.9, applied from iteration 1,000 to 40,000. The training loop was configured to run for a maximum of 40,000 iterations, with validation conducted at intervals of every 4,000 steps. To ensure reproducibility, we set the random seed to 42 and used the AdamW optimizer [58] with an initial learning rate of $1\times 10^{-4}$ and a weight decay of 0.05. The batch size during training was set to 4, and parameters were initialized using the kaiming normal [59]. The segmentation head used in this study was the widely-used Mask2Former [60] to generate masks.
@@ IV-B2 Evaluation Metrics
The metrics mAcc, mIoU, and mDice represent the mean values computed for each class, offering insight into the model’s performance at the class level. In contrast, the overall accuracy (aAcc) evaluates the model’s classification performance across the entire dataset. These metrics are all positive indicators, with higher values signifying better performance, simplifying interpretation, and facilitating model comparison.
@@ IV-C Ablation Studies and Analysis
This section presents a series of comprehensive ablation studies to evaluate the effectiveness of key components in our proposed model using the CloudSEN12_High_L1C dataset.
@@ IV-C1 Variants of VFM’s Backbone
To evaluate the impact of different vision foundation models and their variants on our method’s performance, we conducted comprehensive experiments using SAM [22] and DINOv2 [23] as backbones. Table II compares three available model variants.
We observed that SAM and DINOv2 exhibited consistent performance improvements as the model scale increased. Among the SAM variants, the Huge version performed best, achieving 73.70% mIoU, 84.37% mAcc, 89.95% aAcc, and 84.07% mDice. In contrast, DINOv2 (Large) demonstrated superior performance across all metrics, reaching 74.18% mIoU, 84.79% mAcc, 90.19% aAcc, and 84.46% mDice.
@@
These results indicate that DINOv2 (Large) outperforms SAM (Huge) across all metrics while being significantly more efficient regarding training time and memory usage. Therefore, DINOv2 (Large) emerges as the optimal backbone for our method, striking an excellent balance between performance and computational efficiency for cloud segmentation.
@@ IV-C3 Dimension of Interactive Context
The results show that the model’s performance does not necessarily improve with larger dimensions. Starting from dimension 16 with a mIoU of 73.72%, we observe a slight improvement to 73.77% when increasing to dimension 32. The peak performance is achieved at dimension 64 with a mIoU of 74.18%. However, further increasing the dimension to 128 leads to slight performance degradation (0.44% dropping).
@@ IV-C4 Low-Rank MLP of Adapting Module
The experimental results reveal an interesting relationship between rank size and model performance. With a rank of 4, the model achieves a mIoU of 73.84% while requiring only 0.61M trainable parameters. Increasing the rank to 8 improves performance to 74.03% with 1.02M parameters. The peak performance is reached at rank 16 with a mIoU of 74.18%, which requires 1.82M parameters—further increasing the rank to 32 leads to a slight performance degradation. Based on these observations, we choose rank 16 as our optimal setting. This finding aligns with LoRA’s design principle [46] of achieving efficient adaptation through low-rank approximation.
@@ IV-C5 Importance of the Aggregator
The “Aggregated” strategy, which fuses multiple scale features into a single output with parameter-free operation, demonstrates the best overall performance. It achieves 74.18% mIoU, 84.79% mAcc, and 84.46% mDice while maintaining the same aAcc (90.19%) as the “Single-Scale” approach. This superior performance can be attributed to the strategy’s ability to combine and utilize information from different feature scales effectively. Based on these results, we adopt the “Aggregated” feature output strategy in our final model design, as it provides the most effective use of dense multi-scale feature information without introducing additional parameters.
@@ IV-C6 Interaction Frequency of Adapting Module
The results indicate that the optimal performance is achieved when $N$ is set to 24, where the model attains 74.18% mIoU, 84.79% mAcc, 90.19% aAcc, and 84.46% mDice. This configuration outperforms all others, particularly in terms of mIoU and mAcc. Notably, the increase in interaction frequency from 4 to 24 improves the model’s ability to capture complex spatial dependencies, significantly boosting the performance metrics.